OpenStudy (anonymous):

solve for b: 4(2b-3)=2[3+(b-4)-5b]+2b-6

OpenStudy (dumbcow):

b = 2/7

OpenStudy (anonymous):

this is what I got= 14b=4 b=4/14 or 2/7

OpenStudy (dumbcow):

good job

OpenStudy (anonymous):

b=2/7 use wolfram alpha, it will show you steps and give solution.

OpenStudy (anonymous):

wolfram alpha?